OpenStudy (anonymous):

What is the arc length of a circle that has a 6-inch radius and a central angle that is 65 degrees? Use 3.14 for π and round your answer to the nearest hundredth.

OpenStudy (anonymous):

0.65 inch 1.13 inches 6.80 inches 390.01 inches

OpenStudy (blackstreet23):

formula for arc length \[\int\limits_{a}^{b} \sqrt{1+(f'(x))^2}\]

OpenStudy (anonymous):

so just solve it??

OpenStudy (blackstreet23):

ohh i see

OpenStudy (blackstreet23):

i was having the wrong approach

OpenStudy (blackstreet23):

it is something like this http://www.regentsprep.org/regents/math/algtrig/atm1/arclengthlesson.htm

OpenStudy (blackstreet23):



OpenStudy (blackstreet23):

theta = r / s where s is the arc length

OpenStudy (blackstreet23):

since we want s then our equation is (theta * r)

OpenStudy (blackstreet23):

so it would be (65 * 6)

OpenStudy (blackstreet23):

but that is degrees and we want it in radians

OpenStudy (blackstreet23):

so we need to convert we do that by (65* pi)/180

OpenStudy (blackstreet23):

since they tell you use 3.14 for pi just do (65*3.14)/180

OpenStudy (blackstreet23):

that is approximately 1.13

OpenStudy (blackstreet23):

so plugin back

OpenStudy (blackstreet23):

1.13 * 6

OpenStudy (blackstreet23):

the answer is 6.80

OpenStudy (blackstreet23):

rounding it up

OpenStudy (blackstreet23):

do you get why?

OpenStudy (blackstreet23):

ok good luck then!

OpenStudy (anonymous):

sorry i left my computer i understand it now thank you so much @blackstreet23